OpenStudy (anonymous):

The pair of points is on the graph of an inverse variation. Find the missing value. (2.4, 3) and (5, y)

OpenStudy (zale101):

Inverse Variation: y = k/x

OpenStudy (zale101):

Use K for (2.4, 3)

OpenStudy (zale101):

3 = k/2.4 Solve?

OpenStudy (zale101):

Let's solve the first point and by doing so we have to find the constant of variation which y = k/x and 3 = k/2.4. Using this same constant of variation with the second point, 2.4 *3=7.2 y = k/x y = 7.2/5 y = ?